Observer1, Nicolas Bouche, Peter Buschkamp Observer2, Giacomo Beccari

Inst. Support, Dave Thompson TOs; David Gonzalez, Aaron Ceranski

Nice sunset (means lots of high cirrus clouds), clearly not photometric night, as a front is moving in.

UT 01:15: dome opens

UT 01:30: tel elevation encoder does not want to be reseted. (was worked on during the day due to last night problems)

UT 02:00: more pblms with elevation encoder;

UT 03:45: decision not to use the troubled encoder

UT 04:10: Sig Or OB4 repeat from last night seeing 1.2"

UT 04:50: M67_B1 seeing 1.1 to 1.4

UT 05:35 seeing >1.5"

UT 05:40 M81BV x 2

UT 06:15 clouds rolling

UT 06:45 SN2006jc

UT 07:00 wave front sensor hung

UT 08:15 WFS still stuck; thick clouds

-- NicolasBouche - 18 Feb 2007
Topic revision: r5 - 18 Feb 2007, NicolasBouche
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback